OpenStudy (anonymous):

how much do i need to deposit to get 25000 in 6years 5%compounded semiannually, what is the formula thank you in advance

OpenStudy (anonymous):

Well, compounding semiannually your effective interest rate is defined as: r = (1 + 0.05/2)^2 - 1 = 5.0625% interest, yearly. So over 6 years that's going to be d*(1.050625)^6 where d = deposited amount. So you have d*(1.050625)^6 - d = 25000. Solve for d.

OpenStudy (anonymous):

ty rav will now have a go at it

OpenStudy (anonymous):

A = Final Amount P = Principle (starting balance) r = Interest rate (as a decimal) n = number of compounds per year t = time (in years) \[a=p(1+\frac{ r }{ n})^{nt}\]

OpenStudy (anonymous):

ty will try it now